OpenStudy (anonymous):

Mind to help Checking my Answer? Thanks~ A population of a rare insect in an area will grow at a rate that is proportional to their current population. In the absence of any outside factors, the population will triple in two weeks’ time. On any given day there is a net migration (into the area) of 15 insects and there are also an extinction cause by identified factor such as 16 are eaten by the local bird population and 7 die of natural causes. If there are initially 100 insects in the area, predict either the population of the rare insect will survive or not? If not, when do they will die out

OpenStudy (anonymous):

is it right for me to state that \[\frac{ dx }{ dt }=x(\frac{ 3 }{ 14 })-8 ?\]

OpenStudy (anonymous):

I then get \[dt=\frac{ 14 }{ 3 } ( \frac{ dx }{ x - \frac{ 112 }{ 3 } } )\]\[t + c= \frac{ 14 }{ 3 } \ln (x - \frac{ 112 }{ 3 } )\]\[\frac{ 3 }{ 14 }t + C= \ln(x-\frac{ 112 }{ 3 } )\]\[x - \frac{ 112 }{ 3 } = e^{\frac{ 3 }{ 14 }} C\]After that I sub the given value and get C, and sub x=0 to get t. But I'm wondering, am I correct?

OpenStudy (anonymous):

I'm not sure am I right about the rate of increasing, which I put as 3/14

OpenStudy (loser66):

I don't know how to construct the equation, but I know how to solve it.

OpenStudy (anonymous):

please help me :)

OpenStudy (twopointinfinity):

If I give you an invisible cupcake will you give me a medal?

OpenStudy (anonymous):

I'm willing to give medal to someone who can help me out :)

OpenStudy (anonymous):

Must your approach involve a differential equation? I notice that this is inherently a discrete situation and that the numbers involved are small. I can choose a rate of increase so that the population triples in 14 steps: 0 107.597 1 115.77114409 2 124.566277907 3 134.029578039 4 144.211805083 5 155.167575915 6 166.955656657 7 179.639277893 8 193.286473835 9 207.970447252 10 223.76996213 11 240.769766153 12 259.061045288 13 278.741912898 14 299.917936021 Then I can decrement the increase at each step by 8 to see what effect that has on the population size. 0 99.597 1 99.16338409 2 98.6968263793 3 98.1948242794 4 97.6546850799 5 97.0735115054 6 96.4481861744 7 95.7753548781 8 95.0514085882 9 94.2724640986 10 93.4343431962 11 92.5325502488 12 91.5622480912 13 90.5182320787 14 89.3949021698 Clearly the population is dying out.

OpenStudy (anonymous):

Thanks for the respond, so the population is an exponential growth? I thought the absence of any outer factor will make it a directly proportional growth, which is 3/14

OpenStudy (nincompoop):

what is your C?

OpenStudy (anonymous):

I dint calculate out the C, but I think we can get it by substitute x= 100 and t=0?

OpenStudy (anonymous):

The question says that population growth rate is proportional to population size, and that was reflected in your original work. As I see it, the difficulty with this problem is that counts are taken in discrete time yet you might prefer to use differential equations which are essentially continuous.

OpenStudy (anonymous):

I think I can see the picture already, thanks so much :)

OpenStudy (anonymous):

You're most welcome.